If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by checkley71(8403)

5/20+10/20+15/20=.25+.5+.75 AVERAGE TIME FOR 3 DRINKS IS .5 MIN. THUS FOR 20 DRINKS WE GET 3/20=.5/X OR 3X=10 OR X=3 1/3 MIN.